OpenStudy (anonymous):

so this one would be set up how? Suppose the area of a rectangle is 22.7 ft 2 and the length is 11 ft longer than the width. Find the length and width in feet of the rectangle.

OpenStudy (anonymous):

let width =x length = x+11

OpenStudy (anonymous):

x(x+11) = 22.7 x^2 + 11x - 22.7 = 0

OpenStudy (anonymous):

solve and use positive value

OpenStudy (anonymous):

width is 1.78 length is 12.78